6).

a.The accompanying data represent the miles per gallon of a random sample of cars with a three-cylinder, 1.0 liter engine. (a)b.Twenty girls (ages 9-10) competed in the 50-meter freestyle event at a local swim meet. The mean time was 43.70 seconds with
Answer #1

a)Z score=(x-mean)/s.d

mean of the MPG=39.025

standard deviation of the MPG=3.49

therefore z score of 43.7 = 1.33

The z score is 1.33 standard deviations avove the mean

b) Since there are even no. of data, the quartiles are

Q1=(n/4 th value + (n/4 +1)th value)/2 = 36.55

Q2=(n/2 th value + (n/2 +1)th value)/2=38.55

Q3=(3n/4 th value + (3n/4 +1)th value)/2=41.375

c)The IQR= Q3-Q1=4.825

The IQR is 4.825. It is the range of the middle 50% observations in the data set

d)The lower fence=Q1-1.5*IQR = 29.31

The upper fence=Q3+1.5*IQR= 48.61

since all the data points are between the lower and upper fences there are no outliers

6b)option C. The distribution is likely to be skewed to the right since the mean is greater than the median. The presence of outliers on the high end could explain the higher standard deviation and the greater mean.